OpenStudy (anonymous):

The single equivalent discount rate of the trade discount 5/4/1 is:

OpenStudy (dumbcow):

what do you mean by "5/4/1"

OpenStudy (anonymous):

That is the way the question was posed to me. I am not sure I understand the question either.

OpenStudy (anonymous):

If there are 3 successive discount, x, y and z then net discount should be the following: = (x + y + z) – [(xy/100) + (yz/100) + (xz/100) – (xyz/10000)] = (5 + 4 +1) – [(20/100) + (4/100) + (5/100) – (20/10000)] = 9.71 % = 0.0971..... there ya go =)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Thank you!